OpenStudy (anonymous):

4-6|3-5| divided by 3

OpenStudy (anonymous):

0

OpenStudy (anonymous):

4-6=-2 |3-5| means the absolute value of 3-5, so that would be 2, -2 * 2 = -4/3 = -1.3 Is that a choice? o,o

OpenStudy (anonymous):

\[\left| 3-5 \right|\] \[4-12=-8\] \[-8/3\]

OpenStudy (anonymous):

\[\left| 3-5 \right|=2\]